OpenStudy (anonymous):

If g(x) = 2(x - 4), find the value of x if g(x) = 20

OpenStudy (whpalmer4):

If g(x) = 2(x-4) and g(x) = 20 20 = 2(x-4) Solve for x

OpenStudy (nikato):

Substitute 20=2(x-4) Distribute 20=2x-8 Add 8 to both sides 28=2x Divide both sides by 2 x=14

OpenStudy (whpalmer4):

Or: 20=2(x-4) Divide by 2 10 = x-4 add 4 14=x